# ControlAudio: Tackling Text-Guided, Timing-Indicated and Intelligible Audio Generation via Progressive Diffusion Modeling

### v1

@@ A.2 Timing-Controlled Datasets
For the timing-control fine-tuning stage, our dataset is constructed based on AudioSet-Strong Hershey et al. (2021), which contains 1.8M audio clips. This dataset is crucial as it provides dense, frame-level timestamps for 456 sound event classes. However, since AudioSet-Strong only provides categorical labels (e.g., "Dog"), not descriptive text, we generate richer captions for each timed event. Inspired by the methodology of WavCaps Mei et al. (2024), we employ a large language model (LLM) to create a unique textual description for each segmented audio event.
@@ A.3 Annotated Data Pipeline
Initial Data Selection from AudioSet-SL. We begin with the AudioSet-SL dataset Hershey et al. (2021), which contains strong, human-verified temporal annotations for a wide range of sound events. From the full dataset, we first identify and select all 10-second audio clips that contain at least one event labeled as "Human speech," "Speech," or any of their subcategories. This initial filtering yields a subset of 49,950 clips containing speech.
High-Quality Speech Track Extraction. To obtain high-quality and reliable speech stems, we employ a dual-demixing comparison strategy inspired by MTV Weng et al. (2025). This strategy involves comparing the separation outputs from MVSEP Solovyev and Spleeter Hennequin et al. (2020) to filter for quality and extract a clean speech signal for the subsequent processing steps.
@@ A.4 Simulated Data Pipeline
Guided Synthesis Pipeline. The synthesis process for each 10-second clip is as follows. First, we determine the scenario type by sampling from the speaker distribution (a 79.1% chance of a single-speaker monologue). Next, we source clean speech utterances with transcripts from the LibriTTS-R dataset Koizumi et al. (2023). For a monologue, we sample a number of utterances (determined by the utterance-per-clip distribution, and capped at a maximum of 8) from a single speaker. For a dialogue, we sample utterances from 2 to 4 different speakers, ensuring that no single speaker contributes more than 4 utterances. We then simulate a plausible temporal arrangement for these utterances within the 10-second window. Finally, the composed speech-only track is mixed with a non-speech background audio clip randomly selected from a filtered subset of WavCaps Mei et al. (2024) and VGG-Sound Chen et al. (2020). The mixing is performed at a signal-to-noise ratio (SNR) randomly sampled from a uniform distribution between 2 and 10 dB.
@@ Appendix B Model Configurations
The diffusion network backbone is a DiT with 24 layers, 24 attention heads, and a model hidden dimension of 1536 Evans et al. (2024b). The model utilizes both cross-attention for all conditional inputs and global conditioning for duration-related signals. The internal token dimension of the diffusion model is 64, with a conditional token dimension of 768 and a global condition embedding dimension of 1536.
@@ B.1 Compression Networks
Our audio autoencoder is a variational autoencoder (VAE) based on the Descript Audio VAE Evans et al. (2025) framework, operating at a 16kHz sampling rate. The model is trained from scratch on the audio portions of large-scale public datasets to learn a compact audio representation. The encoder is configured with a model dimension (d_model) of 128 and uses strides of [4, 4, 4, 10], resulting in an overall downsampling ratio of 640. The encoder maps the input waveform into a final 64-dimensional latent representation, which is then used by the decoder for reconstruction. The model’s input/output channels (io_channels) are set to 1 for mono audio. We use Snake activation throughout the network and omit the final tanh activation in the decoder.
@@ B.2 Training Details
To improve convergence stability and generation quality, we adopt several common training strategies Evans et al. (2025), with configurations specified in our training setup. We apply Exponential Moving Average (EMA) to the model parameters. For optimization, we use the AdamW optimizer with a learning rate of $5\times 10^{-5}$, $(\beta_{1},\beta_{2})=(0.9,0.999)$, and a weight decay of $1\times 10^{-3}$.
@@ C.1 Objective Metrics
Audio Quality. Our primary metric for audio fidelity is the Fréchet Audio Distance (FAD), which measures the distributional difference between generated and reference audio based on VGGish embeddings. To evaluate the consistency of acoustic event distributions, we also report Kullback-Leibler (KL) divergence computed using the PANNs tagging model. For completeness and comparison with prior works, we include the Inception Score (IS) and Fréchet Distance (FD) as supplementary metrics Liu et al. (2023).
Semantic Alignment. To measure the alignment between the generated audio and its corresponding text prompt, we use the LAION-CLAP Wu et al. (2023) score. This score is defined as the cosine similarity between the CLAP embeddings of the generated audio $a$ and the text prompt $t$:
@@ D.1 Chain-of-Thought for Prompt Planning
Recent advances have demonstrated the powerful planning and cross-modal reasoning capabilities of large language models (LLMs) Wang et al. (2025b, 2024). We leverage these capabilities by employing an LLM to function as a "planner" that automatically converts a free-form natural language caption ($y_{c}$) into a precise, structured prompt ($y_{s}$) for our generative model. This conversion follows a three-stage reasoning process inspired by the Chain-of-Thought (CoT) paradigm, as illustrated in Figure 4. The process consists of the following steps:

### v2

@@ A.2 Timing-Controlled Datasets
For the timing-control fine-tuning stage, our dataset is constructed based on AudioSet-Strong hershey2021benefit, which contains 1.8M audio clips. This dataset is crucial as it provides dense, frame-level timestamps for 456 sound event classes. However, since AudioSet-Strong only provides categorical labels (e.g., "Dog"), not descriptive text, we generate richer captions for each timed event. Inspired by the methodology of WavCaps mei2024wavcaps, we employ a large language model (LLM) to create a unique textual description for each segmented audio event.
@@ A.3 Annotated Data Pipeline
Initial Data Selection from AudioSet-SL. We begin with the AudioSet-SL dataset hershey2021benefit, which contains strong, human-verified temporal annotations for a wide range of sound events. From the full dataset, we first identify and select all 10-second audio clips that contain at least one event labeled as "Human speech," "Speech," or any of their subcategories. This initial filtering yields a subset of 49,950 clips containing speech.
High-Quality Speech Track Extraction. To obtain high-quality and reliable speech stems, we employ a dual-demixing comparison strategy inspired by MTV weng2025audio. This strategy involves comparing the separation outputs from MVSEP ZFTurbo23 and Spleeter hennequin2020spleeter to filter for quality and extract a clean speech signal for the subsequent processing steps.
@@ A.4 Simulated Data Pipeline
Guided Synthesis Pipeline. The synthesis process for each 10-second clip is as follows. First, we determine the scenario type by sampling from the speaker distribution (a 79.1% chance of a single-speaker monologue). Next, we source clean speech utterances with transcripts from the LibriTTS-R dataset koizumi2023libritts. For a monologue, we sample a number of utterances (determined by the utterance-per-clip distribution, and capped at a maximum of 8) from a single speaker. For a dialogue, we sample utterances from 2 to 4 different speakers, ensuring that no single speaker contributes more than 4 utterances. We then simulate a plausible temporal arrangement for these utterances within the 10-second window. Finally, the composed speech-only track is mixed with a non-speech background audio clip randomly selected from a filtered subset of WavCaps mei2024wavcaps and VGG-Sound chen2020vggsound. The mixing is performed at a signal-to-noise ratio (SNR) randomly sampled from a uniform distribution between 2 and 10 dB.
@@ Appendix B Model Configurations
The diffusion network backbone is a DiT with 24 layers, 24 attention heads, and a model hidden dimension of 1536 evans2024long. The model utilizes both cross-attention for all conditional inputs and global conditioning for duration-related signals. The internal token dimension of the diffusion model is 64, with a conditional token dimension of 768 and a global condition embedding dimension of 1536.
@@ B.1 Compression Networks
Our audio autoencoder is a variational autoencoder (VAE) based on the Descript Audio VAE evans2025stable framework, operating at a 16kHz sampling rate. The model is trained from scratch on the audio portions of large-scale public datasets to learn a compact audio representation. The encoder is configured with a model dimension (d_model) of 128 and uses strides of [4, 4, 4, 10], resulting in an overall downsampling ratio of 640. The encoder maps the input waveform into a final 64-dimensional latent representation, which is then used by the decoder for reconstruction. The model’s input/output channels (io_channels) are set to 1 for mono audio. We use Snake activation throughout the network and omit the final tanh activation in the decoder.
@@ B.2 Training Details
To improve convergence stability and generation quality, we adopt several common training strategies evans2025stable, with configurations specified in our training setup. We apply Exponential Moving Average (EMA) to the model parameters. For optimization, we use the AdamW optimizer with a learning rate of $5\times 10^{-5}$, $(\beta_{1},\beta_{2})=(0.9,0.999)$, and a weight decay of $1\times 10^{-3}$.
@@ C.1 Objective Metrics
Audio Quality. Our primary metric for audio fidelity is the Fréchet Audio Distance (FAD), which measures the distributional difference between generated and reference audio based on VGGish embeddings. To evaluate the consistency of acoustic event distributions, we also report Kullback-Leibler (KL) divergence computed using the PANNs tagging model. For completeness and comparison with prior works, we include the Inception Score (IS) and Fréchet Distance (FD) as supplementary metrics liu2023audioldm.
Semantic Alignment. To measure the alignment between the generated audio and its corresponding text prompt, we use the LAION-CLAP wu2023large score. This score is defined as the cosine similarity between the CLAP embeddings of the generated audio $a$ and the text prompt $t$:
@@ D.1 Chain-of-Thought for Prompt Planning
Recent advances have demonstrated the powerful planning and cross-modal reasoning capabilities of large language models (LLMs) wang2025multimodal; wang2024qwen2. We leverage these capabilities by employing an LLM to function as a "planner" that automatically converts a free-form natural language caption ($y_{c}$) into a precise, structured prompt ($y_{s}$) for our generative model. This conversion follows a three-stage reasoning process inspired by the Chain-of-Thought (CoT) paradigm, as illustrated in Figure 4. The process consists of the following steps: